OpenStudy (tad1):

I posted this one a few days ago, but no oe answered it. Please try. Consider the table 1 = 0 + 1 1 +3 +4 = 1 + 8 5 + 6 + 7+8+9 = 8 + 27 10+11+12+13+14+15+16 = 27 +64 each line has the next odd number of consecutive integers, and each line equals the sum of the last perfect cube used and the next consecutive cube. but I can't figure out how to write this in suitable mathematical form, let alone prove it.

OpenStudy (anonymous):

this looks like an identity of some sort but i notice you have omitted the number 2

OpenStudy (anonymous):

oh the second line is 2 + 3+ 4, right?

OpenStudy (anonymous):

i think 2 was there in the second line becoz 2 + 3 + 4 = 1+8

OpenStudy (tad1):

yes, sorry for the typo

OpenStudy (anonymous):

they are all arithmetic series with d = 1 and different first term

OpenStudy (anonymous):

but where do we go from there?

OpenStudy (tad1):

but how would you write it in mathematical form so that you could proove what the next line would be?

OpenStudy (anonymous):

hmm - good question!!

OpenStudy (tad1):

This is problem #2 in Chapter one in George Polya"s Introduction and Analogy in Mathematics,Vol 1. I think it's supposed to be solved with induction. Can we do induction on a list of arithmetic series?

OpenStudy (tad1):

title of book is Induction and Analogy

OpenStudy (anonymous):

here's some bits:- sum of first k digits k(k+1)/2 and sum of first k cubes is /k(k+1)/2)^2

OpenStudy (anonymous):

i expect we can

OpenStudy (tad1):

I like where we're going. But how do we indicate that the next line starts at the next consecutive number each time, and that the cubes are also consecutive?

OpenStudy (tad1):

i don't think you expression for cubes works 0 + 1)/2 =0 1^3 + 2^3)/2 =3

OpenStudy (anonymous):

i'm not ducking out of this tad but I have to play taxi driver to my grandaughter i'll be back later

OpenStudy (tad1):

= 4.5 not 3

OpenStudy (tad1):

thanks jimmy

OpenStudy (anonymous):

(k(k+1)/2)^2 1(1+1)/2 squared = 1 2(2+1)/2 squared = 9

OpenStudy (tad1):

but the first cube is 0

OpenStudy (anonymous):

Number theory doesn't consider 0

OpenStudy (tad1):

this problem does

OpenStudy (anonymous):

Just start your induction at 1...

OpenStudy (tad1):

okay; if I do that, how do I indicate that each line has different numbers, and a different number of numbers

OpenStudy (anonymous):

The last number in a line n on the lhs is n^2 and the rest of the numbers are consecutive down to (n-1)^2 +1 The rhs for a line n is 1 + 2^3 + 3^3 +....n^3

OpenStudy (tad1):

that's cool. I think i can get some where. Be back in a while